# Sharpness of Minima in Deep Matrix Factorization: Exact Expressions

### v4

@@ Abstract
Understanding the geometry of the loss landscape near a minimum is key to explaining the implicit bias of gradient-based methods in non-convex optimization problems such as deep neural network training and deep matrix factorization. A central quantity to characterize this geometry is the maximum eigenvalue of the Hessian of the loss, which measures the sharpness of the landscape. Currently, its precise role has been obfuscated because no exact expressions for this sharpness measure were known in general settings. In this paper, we present the first exact expression for the maximum eigenvalue of the Hessian of the squared-error loss at any minimizer in general overparameterized deep matrix factorization (i.e., deep linear neural network training) problems, resolving an open question posed by Mulayoff & Michaeli (2020). This expression uncovers a fundamental property of the loss landscape of depth-2 matrix factorization problems: a minimum is flat if and only if it is spectral-norm balanced, which implies that flat minima are not necessarily Frobenius-norm balanced. Furthermore, to complement our theory, we empirically investigate an escape phenomenon observed during gradient-based training near a minimum that crucially relies on our exact expression of the sharpness.
@@ 1 Introduction
Decades of research in learning theory suggest limiting model complexity to prevent overfitting. However, modern deep learning is heavily overparameterized and has nonetheless achieved unprecedented success in practice over the past decade (Krizhevsky et al., 2012; Vaswani et al., 2017). Generally, in overparameterized settings, the loss function has infinitely many global minima that achieve zero training error (interpolation regime), yet these models still perform well. This phenomenon has been explored in various settings such as nonparametric regression, (Belkin et al., 2019), training two-layer neural networks with logistic loss (Frei et al., 2022), and linear regression (Bartlett et al., 2020).
The propensity of neural network training dynamics to converge to good minima is attributed to the ability of gradient-based optimization algorithms to avoid bad minima (Neyshabur et al., 2017; Zhang et al., 2017). This is related to the implicit bias of gradient descent (GD) (Neyshabur et al., 2014), and a large body of work has focused on its understanding (Gunasekar et al., 2017; 2018; Soudry et al., 2018; Arora et al., 2019; Ji & Telgarsky, 2020; Yun et al., 2021).
It has been observed that dynamical stability of GD near a minimum is a key factor in characterizing its implicit bias toward particular solutions (Wu et al., 2018; Nar & Sastry, 2018). Conceptually, dynamical stability refers to the ability of GD to stably converge to a minimum, and it is closely related to the sharpness of the loss landscape in its vicinity (Mulayoff et al., 2021). This topic has been investigated in numerous works (Nar & Sastry, 2018; Wu et al., 2018; Ma & Ying, 2021; Mulayoff et al., 2021; Nacson et al., 2023; Qiao et al., 2024; Liang et al., 2025) within the framework of the classical notion of linear stability in dynamical systems (Strogatz, 2024).
@@
Table 1: Closed-form expressions for the maximum Hessian eigenvalue in the literature. $\Omega$ denotes the set of all global minimizers, $\Omega_{F}\subseteq\Omega$ denotes the set of flat global minimizers, and $\Omega_{B}\subseteq\Omega$ denotes the set of balanced global minimizers.
Related Work Depth Input Dim. Output Dim. $\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}))$ Layers
Mulayoff & Michaeli (2020, Theorem 1) $L$ $d_{0}$ $d_{L}$ $\bm{w}\in\Omega_{F}$ $\mathbb{R}^{a\times b}$
Zhu et al. (2023, Appendix B.1) 22 11 11 $\bm{w}\in\mathbb{R}^{N}$ $\mathbb{R}$
Singh & Hofmann (2024, Theorem 1) 22 11 11 $\bm{w}\in\mathbb{R}^{N}$ $\mathbb{R}^{a}$
Ghosh et al. (2025, Lemma 1) $L$ $d_{0}$ $d_{L}$ $\bm{w}\in\Omega_{B}$ $\mathbb{R}^{a\times b}$
Theorem 5 (This Paper) $L$ $d_{0}$ $d_{L}$ $\bm{w}\in\Omega$ $\mathbb{R}^{a\times b}$
Most notably, the seminal work of Mulayoff & Michaeli (2020) derives a closed-form expression for the maximum eigenvalue of the Hessian at flat global minima of deep linear networks (i.e., deep matrix factorization) with squared-error loss. However, obtaining a closed-form expression for all global minima in deep linear networks/deep matrix factorization was an open problem. In particular, Mulayoff & Michaeli (2020) claim that finding a closed-form expression for arbitrary global minima is intractable. In this paper, we refute this claim and positively answer the following fundamental question.
@@
In particular, in Theorem 5, we provide a closed-form expression for the maximum Hessian eigenvalue at arbitrary minima of depth-$L$ overparameterized deep matrix factorization. We also highlight that overparameterized deep matrix factorization and the deep linear neural network setting of Mulayoff & Michaeli (2020) are equivalent for investigating the sharpness measure $\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}^{*}))$ when the data covariance matrix is identity, i.e., $\bm{\Sigma}_{x}=\bm{I}$ (see Mulayoff & Michaeli, 2020, Equation 21). To the best of our knowledge, our analysis provides the first exact expression of the maximum eigenvalue for deep matrix factorization/deep linear neural network problems. In the case of deep overparameterized scalar factorization (Theorem 21) and depth-$2$ matrix factorization (Corollary 25), our closed-form expression simplifies considerably.
With our closed-form expression in hand, we then empirically explore in Section 7 the escape phenomenon, observed by Wu et al. (2018) (who only studied a one-dimensional setting). We find that this phenomenon also occurs for overparameterized deep matrix factorization problems. Therefore, we empirically observe the following.
GD escapes from a dynamically unstable minimum almost surely.
We explore this phenomenon through the lens of dynamical stability introduced by Wu et al. (2018).
@@
Consider a twice continuously differentiable loss function $\mathcal{L}:\mathbb{R}^{N}\rightarrow\mathbb{R}$, and let $\bm{x}^{*}$ be a minimizer of $\mathcal{L}$. For $\delta>0$, denote the open ball $B_{\delta}(\bm{x}^{*}):=\left\{\bm{x}\in\mathbb{R}^{N}:\norm{\bm{x}-\bm{x}^{*}}_{2}<\delta\right\}$. There exists a $\delta>0$ such that, for all $\bm{x}\in B_{\delta}(\bm{x}^{*})$,
$\mathcal{L}(\bm{x})=\mathcal{L}(\bm{x}^{*})+(\bm{x}-\bm{x}^{*})^{\top}\nabla\mathcal{L}(\bm{x}^{*})+\frac{1}{2}(\bm{x}-\bm{x}^{*})^{\top}\nabla^{2}\mathcal{L}(\bm{x}^{*})(\bm{x}-\bm{x}^{*})+o(\norm{\bm{x}-\bm{x}^{*}}_{2}^{2}).$ (1)
Definition 1.
@@
Let $\mathcal{L}:\mathbb{R}^{N}\rightarrow\mathbb{R}$ be a twice continuously differentiable loss function, and let $\bm{x}^{*}$ be a minimizer of $\mathcal{L}$. The corresponding linearized GD update rule for $\mathcal{L}$ is given by
$\bm{x}_{t+1}=\bm{x}_{t}-\eta\nabla^{2}\mathcal{L}(\bm{x}^{*})(\bm{x}_{t}-\bm{x}^{*}),$ (2)
@@
where $\eta>0$ is the step size. Given $\delta>0$ such that (1) holds, if there exists an initial point $\bm{x}_{0}\in B_{\delta}(\bm{x}^{*})$ such that the residuals $\bm{\epsilon}_{t}=\bm{x}_{t}-\bm{x}^{*}$ diverge, i.e.,
@@
then $\bm{x}^{*}$ is said to be dynamically unstable.
A necessary and sufficient condition for a minimum to be dynamically unstable is that $\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{x}^{*}))>2/\eta$ (cf. Wu et al. (2018); Mulayoff et al. (2021); Chemnitz & Engel (2025)). Thus, we see that an exact expression for the maximum eigenvalue is necessary to explore the escape phenomenon empirically.
@@ 1.1 Contributions
  • A minimizer of deep overparameterized scalar factorization loss is flat if and only if the product of spectral norms of left and right intermediate factors is constant across layers (Corollary 7).
  • Flat minima are spectral-norm balanced in depth-2 matrix factorization (Corollary 33). This implies that flat minima are not necessarily balanced, contrary to claims made in several works (Ding et al., 2024; Ghosh et al., 2025). We further discuss this in Section 6.
  • A minimizer of deep matrix factorization loss is flat if the product of spectral norms of left and right intermediate factors is constant across layers (Corollary 10).
  • A recent work (Anonymous, 2025) provides empirical evidence that, in two-layer linear networks, the largest eigenvalue of the Hessian of the loss function is correlated with the spectral norm of the weight matrices. Furthermore, for deep linear networks, they observe that this largest Hessian eigenvalue correlates with the spectral norm of the products of the weight matrices. Our results in Theorem 21, Theorem 5, and Corollary 25 provide the first theoretical explanation for their empirical observations.
@@ 1.2 Related Work
Flat Minima. Mulayoff & Michaeli (2020) derived a closed-form expression for the maximum eigenvalue of the Hessian at flat minima for deep linear neural networks. They also showed that the Hessian at a global minimum is rank-deficient by at least the order of 11/L1-1/L, where $L$ is the depth of the network. Moreover, they showed that the sharpness of the flattest minima increases approximately linearly with $L$ if $L\gg 1$. Singh & Hofmann (2024) provided a full characterization of the Hessian spectrum at a point in parameter space for linear and ReLU networks in the scalar regression case. They observed that the eigenvalues scale in proportion to the input variance within one hidden-layer scalar linear networks. More recently, Josz (2025) has shown that locally flat minima are globally flat in depth-$2$ matrix factorization problems.
Balanced Minima. Ghosh et al. (2025) provided a full characterization of the Hessian spectrum at balanced minima in deep matrix factorization. Furthermore, they showed that the maximum eigenvalue of the Hessian at the flattest minima is equal to that of the balanced minima. Ding et al. (2024) showed that norm-minimal, balanced, and flat solutions coincide in depth-$2$ matrix factorization, where sharpness measured by the scaled trace of the Hessian matrix of the loss function. We further discuss Ghosh et al. (2025) and Ding et al. (2024)’s results in Section 6. Finally, Wang et al. (2022) showed that large step size GD training induces a balancing effect between factors in depth-$2$ matrix factorization.
@@
Sharpness and Generalization. It is widely recognized in the literature that flat minima are associated with better generalization (Hochreiter & Schmidhuber, 1997; Keskar et al., 2017). In a large-scale empirical investigation, Jiang et al. (2020) examined different measures for deep networks and found that a sharpness-based measure exhibited the strongest correlation with generalization. There is also theoretical evidence for this phenomenon in low-rank matrix recovery (Ding et al., 2024). On the other hand, Dinh et al. (2017) showed that good minima can be arbitrarily sharp in deep neural networks.
@@ 2 Notation, Preliminaries, and Problem Setup
$\prod_{j=n}^{m}\bm{W}_{j}:=\begin{cases}\bm{W}_{m}\bm{W}_{m-1}\dots\bm{W}_{n}\quad&\text{if $n\leq m$},\\ \bm{I}_{d_{m}}&\text{otherwise},\text{ where }n,m\in[L],\end{cases}$ (4)
where $\bm{W}_{m}\in\mathbb{R}^{d_{m}\times d_{m-1}}$.
@@
Gâteaux Derivatives. Let $f:\mathbb{R}^{K\times L}\to\mathbb{R}$ be a differentiable function with continuous first- and second-order derivatives on $\mathbb{R}^{K\times L}$. Our objective is to derive closed-form expressions for the first- and second-order directional derivatives of $f$ in the direction of $\bm{U}\in\mathbb{R}^{K\times L}$, where $\norm{\bm{U}}_{F}<\infty$, denoted respectively by $D_{\bm{U}}f(\bm{X})$ and $D_{\bm{U}}^{2}f(\bm{X})$. By the limit definition of the derivative, the first derivative of $f(\bm{X})$ with respect to each entry of $\bm{X}$ can be expressed as follows:
$\frac{\partial f(\bm{X})}{\partial X_{ij}}=\lim_{\Delta t\rightarrow 0}\frac{f(\bm{X}+\Delta t\bm{e}_{i}\bm{e}_{j}^{\top})-f(\bm{X})}{\Delta t},\quad\forall(i,j)\in[K]\times[L],$ (5)
where $\bm{e}_{i}$ is the $i^{\text{th}}$ standard basis vector of $\mathbb{R}^{K}$ and $\bm{e}_{j}$ is the $j^{\text{th}}$ standard basis vector of $\mathbb{R}^{L}$. If the limit in (5) exists then by substitution of variables
$\frac{\partial f(\bm{X})}{\partial X_{ij}}U_{ij}=\lim_{\Delta t\rightarrow 0}\frac{f(\bm{X}+\Delta tU_{ij}\bm{e}_{i}\bm{e}_{j}^{\top})-f(\bm{X})}{\Delta t},\quad\forall(i,j)\in[K]\times[L].$ (6)
By definition, the total change in $f(\bm{X})$ in the direction of $\bm{U}$ is the sum of change due to each entry of $\bm{X}$. Then
$D_{\bm{U}}f(\bm{X})$ $=\sum_{i,j\in[K]\times[L]}\frac{\partial f(\bm{X})}{\partial X_{ij}}U_{ij}$ (7)
$=\sum_{i,j\in[K]\times[L]}\lim_{\Delta t\rightarrow 0}\frac{f(\bm{X}+\Delta tU_{ij}\bm{e}_{i}\bm{e}_{j}^{\top})-f(\bm{X})}{\Delta t}$ (8)
$=\lim_{\Delta t\rightarrow 0}\frac{f(\bm{X}+\Delta t\bm{U})-f(\bm{X})}{\Delta t}.$ (9)

### v5

@@ Abstract
Understanding the geometry of the loss landscape near a minimum is key to explaining the implicit bias of gradient-based methods in non-convex optimization problems such as deep neural network training and deep matrix factorization. A central quantity to characterize this geometry is the maximum eigenvalue of the Hessian of the loss. Currently, its precise role has been obfuscated because no exact expressions for this sharpness measure were known in general settings. In this paper, we present the first exact expression for the maximum eigenvalue of the Hessian of the squared-error loss at any minimizer in deep matrix factorization/deep linear neural network training problems, resolving an open question posed by Mulayoff and Michaeli (2020). This expression reveals a fundamental property of the loss landscape in deep matrix factorization: Having a constant product of the spectral norms of the left and right intermediate factors across layers is a sufficient condition for flatness. Most notably, in both depth-$2$ matrix and deep overparameterized scalar factorization, we show that this condition is both necessary and sufficient for flatness, which implies that flat minima are spectral-norm balanced even though they are not necessarily Frobenius-norm balanced. To complement our theory, we provide the first empirical characterization of an escape phenomenon during gradient-based training near a minimizer of a deep matrix factorization problem.
Machine Learning, ICML
@@ 1 Introduction
Decades of research in learning theory suggest limiting model complexity to prevent overfitting. However, modern deep learning is heavily overparameterized and has nonetheless achieved unprecedented success in practice over the past decade (Krizhevsky et al., 2012; Vaswani et al., 2017).
Generally, in overparameterized settings, the loss function has infinitely many global minima that achieve zero training error (interpolation regime), yet these models still perform well. This phenomenon has been explored in various settings such as nonparametric regression, (Belkin et al., 2019), training two-layer neural networks with logistic loss (Frei et al., 2022), and linear regression (Bartlett et al., 2020).
The observation of this phenomenon naturally raises the question: Why do optimization algorithms find minima that generalize well? In the literature, the propensity of neural network training dynamics to converge to good minima is generally attributed to the ability of gradient-based optimization algorithms to avoid bad minima (Neyshabur et al., 2017; Zhang et al., 2017). This is related to the implicit bias of gradient descent (GD) (Neyshabur et al., 2014), and a large body of work has focused on its understanding (Gunasekar et al., 2017, 2018; Soudry et al., 2018; Arora et al., 2019; Ji and Telgarsky, 2020; Yun et al., 2021).
Table 1: Closed-form expressions for the maximum Hessian eigenvalue in the literature. $\Omega$ denotes the set of all global minimizers, $\Omega_{F}\subseteq\Omega$ denotes the set of flat global minimizers, and $\Omega_{B}\subseteq\Omega$ denotes the set of Frobenius-balanced global minimizers.
Related Work Depth Input Dim. Output Dim. $\lambda_{\max}(\nabla^{2}\mathcal{L}(\mathbf{w}))$ Layers
Mulayoff and Michaeli (2020, Theorem 1) $L$ $d_{0}$ $d_{L}$ $\mathbf{w}\in\Omega_{F}$ $\mathbb{R}^{a\times b}$
Zhu et al. (2023, Appendix B.1) 22 11 11 $\mathbf{w}\in\mathbb{R}^{N}$ $\mathbb{R}$
Singh and Hofmann (2024, Theorem 1) 22 11 11 $\mathbf{w}\in\mathbb{R}^{N}$ $\mathbb{R}^{a}$
Ghosh et al. (2025, Lemma 1) $L$ $d_{0}$ $d_{L}$ $\mathbf{w}\in\Omega_{B}$ $\mathbb{R}^{a\times b}$
Theorem 3.1 (This Paper) $L$ $d_{0}$ $d_{L}$ $\mathbf{w}\in\Omega$ $\mathbb{R}^{a\times b}$
It has been observed that dynamical stability of GD near a minimum is a key factor in characterizing its implicit bias toward particular solutions (Wu et al., 2018; Nar and Sastry, 2018). Conceptually, dynamical stability refers to the ability of GD to stably converge to a minimum, and it is closely related to the sharpness of the loss landscape in its vicinity (Mulayoff et al., 2021). This topic has been investigated in numerous works (Nar and Sastry, 2018; Wu et al., 2018; Ma and Ying, 2021; Mulayoff et al., 2021; Nacson et al., 2023; Qiao et al., 2024; Liang et al., 2025) within the framework of the classical notion of linear stability in dynamical systems (Strogatz, 2024).
@@
Most notably, the seminal work of Mulayoff and Michaeli (2020) derives a closed-form expression for the maximum eigenvalue of the Hessian at flat global minima of deep linear networks (i.e., deep matrix factorization) with squared-error loss. However, obtaining a closed-form expression for all global minima in deep linear networks/deep matrix factorization was an open problem. In particular, Mulayoff and Michaeli (2020) claim that finding a closed-form expression for arbitrary global minima is intractable. In this paper, we refute this claim and positively answer the following fundamental question.
@@
In particular, in Theorem 3.1, we provide a closed-form expression for the maximum Hessian eigenvalue at arbitrary minima of depth-$L$ overparameterized deep matrix factorization. We also highlight that overparameterized deep matrix factorization and the deep linear neural network setting of Mulayoff and Michaeli (2020) are equivalent for investigating the sharpness measure $\lambda_{\max}(\nabla^{2}\mathcal{L}(\mathbf{w}^{*}))$ when the data covariance matrix is identity, i.e., $\mathbf{\Sigma}_{x}=\mathbf{I}$ (see Mulayoff and Michaeli, 2020, Equation 21). Moreover, when the data matrix is full rank, the deep matrix factorization and the deep linear neural network setting of Mulayoff and Michaeli (2020) are equivalent optimization problems (Chou et al., 2024, Equation 4).
To the best of our knowledge, our analysis provides the first exact expression of the maximum eigenvalue for deep matrix factorization/deep linear neural network training problems. In the case of deep overparameterized scalar factorization (Corollary 3.2) and depth-$2$ matrix factorization (Corollary 28), our closed-form expression simplifies considerably. In Section 4, we present new necessary and sufficient conditions for flat minima in various matrix factorization settings. In Section 5, we demonstrate that different notions of sharpness can lead to misleading conclusions, and we illustrate this with an explicit example.
With our closed-form expression in hand, we then empirically observe in Section 6 the escape phenomenon, which is introduced by Wu et al. (2018). This phenomenon has been observed in numerous works; however, it has been characterized empirically only in extremely simplified settings. For instance, Lewkowycz et al. (2020, see Figure 1) observed this phenomenon for a one-hidden layer linear network trained with squared-error loss on a single scalar data point, Cohen et al. (2021) analyzed it for a convex quadratic loss function and most recently, Ghosh et al. (2025, see Figure 4) characterized it empirically for the two-layer scalar factorization loss. Although their analysis is not directly related to the escape phenomenon, Liang and Montúfar (2025) recently studied the convergence regions of GD in deep matrix factorization problems.
In contrast, we provide the first empirical characterization of an escape phenomenon during gradient-based training near a minimizer of a deep matrix factorization problem. We find that the escape phenomenon also occurs for overparameterized deep matrix factorization problems. We explore this phenomenon through the lens of dynamical stability introduced by Wu et al. (2018).
@@
Consider a twice continuously differentiable loss function $\mathcal{L}:\mathbb{R}^{N}\rightarrow\mathbb{R}$, and let $\mathbf{w}^{*}$ be a minimizer of $\mathcal{L}$. For $\delta>0$, denote the open ball $B_{\delta}(\mathbf{w}^{*}):=\left\{\mathbf{w}\in\mathbb{R}^{N}:\norm{\mathbf{w}-\mathbf{w}^{*}}_{2}<\delta\right\}$. There exists a $\delta>0$ such that, for all $\mathbf{w}\in B_{\delta}(\mathbf{w}^{*})$,
$\mathcal{L}(\mathbf{w})$ $=\mathcal{L}(\mathbf{w}^{*})+(\mathbf{w}-\mathbf{w}^{*})^{\top}\nabla\mathcal{L}(\mathbf{w}^{*})$ (1)
$+\frac{1}{2}(\mathbf{w}-\mathbf{w}^{*})^{\top}\nabla^{2}\mathcal{L}(\mathbf{w}^{*})(\mathbf{w}-\mathbf{w}^{*})$ (2)
$+o(\norm{\mathbf{w}-\mathbf{w}^{*}}_{2}^{2}).$ (3)
Definition 1.1.
@@
Let $\mathcal{L}:\mathbb{R}^{N}\rightarrow\mathbb{R}$ be a twice continuously differentiable loss function, and let $\mathbf{w}^{*}$ be a minimizer of $\mathcal{L}$. The linearized GD update rule for $\mathcal{L}$ is given by
$\mathbf{w}_{t+1}=\mathbf{w}_{t}-\eta\nabla^{2}\mathcal{L}(\mathbf{w}^{*})(\mathbf{w}_{t}-\mathbf{w}^{*}),$ (4)
@@
where $\eta>0$ is the step size. Given $\delta>0$ such that (1) holds, if there exists an initial point $\mathbf{w}_{0}\in B_{\delta}(\mathbf{w}^{*})$ such that the residuals $\bm{\epsilon}_{t}=\mathbf{w}_{t}-\mathbf{w}^{*}$ diverge, i.e.,
@@
then $\mathbf{w}^{*}$ is said to be dynamically unstable.
A necessary and sufficient condition for a minimum to be dynamically unstable is that $\lambda_{\max}(\nabla^{2}\mathcal{L}(\mathbf{w}^{*}))>2/\eta$ (cf. Wu et al. 2018; Mulayoff et al. 2021; Chemnitz and Engel 2025). Thus, we see that an exact expression for the maximum eigenvalue is necessary to explore the escape phenomenon empirically.
@@ 1.1 Contributions
  • A minimizer of the deep overparameterized scalar factorization loss is flat if and only if the product of spectral norms of the left and right intermediate factors is constant across layers (Corollary 4.1).
  • Flat minima are spectral-norm balanced in depth-$2$ matrix factorization (Corollary 36). This implies that flat minima are not necessarily Frobenius-norm balanced, contrary to claims made in several works (Ding et al., 2024; Ghosh et al., 2025). We further discuss this in Section 5.
  • A minimizer of the deep matrix factorization loss is flat if the product of spectral norms of the left and right intermediate factors is constant across layers (Corollary 4.4).
We also highlight a recent work of Anonymous (2025) that provides empirical evidence, in two-layer linear networks, that the largest eigenvalue of the Hessian of the loss is correlated with the spectral norm of the weight matrices. Furthermore, for deep linear networks, they observe that this largest Hessian eigenvalue correlates with the spectral norm of the products of the weight matrices. Thus, our results in Theorem 3.1, Corollary 3.2, and Corollary 28 provide the first theoretical explanation for their empirical observations.
@@ 1.2 Related Work
Flat Minima. Mulayoff and Michaeli (2020) derived a closed-form expression for the maximum eigenvalue of the Hessian at flat minima for deep linear neural networks. They also showed that the Hessian at a global minimum is rank-deficient by at least the order of 11/L1-1/L, where $L$ is the depth of the network. Moreover, they showed that the sharpness of the flat minima increases approximately linearly with $L$ if $L\gg 1$. Singh and Hofmann (2024) provided a full characterization of the Hessian spectrum at a point in parameter space for linear and ReLU networks in the scalar regression case. They observed that the eigenvalues scale in proportion to the input variance within one hidden-layer scalar linear networks. More recently, Josz (2025) has shown that locally flat minima are globally flat in depth-$2$ matrix factorization problems.
Balanced Minima. Ghosh et al. (2025) provided a full characterization of the Hessian spectrum at Frobenius-norm balanced minima in deep matrix factorization problems. Furthermore, they showed that the maximum eigenvalue of the Hessian at flat minima is equal to that of the Frobenius-norm balanced minima. Ding et al. (2024) showed that norm-minimal, Frobenius-norm balanced and flat solutions coincide in depth-$2$ matrix factorization problems, where the sharpness is measured by the scaled trace of the Hessian matrix of the loss function. We further discuss the results of Ghosh et al. (2025) and Ding et al. (2024) in Section 5. Finally, Wang et al. (2022) showed that large step size GD training induces a Frobenius-norm balancing effect between factors in depth-$2$ matrix factorization problems.
@@
Sharpness and Generalization. It is widely recognized in the literature that flat minima are associated with better generalization (Hochreiter and Schmidhuber, 1997; Keskar et al., 2017). In a large-scale empirical investigation, Jiang et al. (2020) examined different complexity measures for deep networks and found that a sharpness-based measure exhibited the strongest correlation with generalization. There is also theoretical evidence for this phenomenon in low-rank matrix recovery (Ding et al., 2024). On the other hand, Dinh et al. (2017) showed that good minima can be arbitrarily sharp in deep neural networks.
@@ 2 Notation, Preliminaries, and Problem Setup
$\prod_{j=n}^{m}\mathbf{W}_{j}:=\begin{cases}\mathbf{W}_{m}\mathbf{W}_{m-1}\dots\mathbf{W}_{n}\quad&\text{if $n\leq m$},\\ \mathbf{I}_{d_{m}}&\text{o.w.},\forall n,m\in[L],\end{cases}$ (6)
where $\mathbf{W}_{m}\in\mathbb{R}^{d_{m}\times d_{m-1}}$.
@@
2.1 Gâteaux Derivatives
Let $f:\mathbb{R}^{K\times L}\to\mathbb{R}$ be a differentiable function with continuous first- and second-order derivatives on $\mathbb{R}^{K\times L}$. Our objective is to derive closed-form expressions for the first- and second-order directional derivatives of $f$ in the direction of $\mathbf{U}\in\mathbb{R}^{K\times L}$, where $\norm{\mathbf{U}}_{F}<\infty$, denoted respectively by $D_{\mathbf{U}}f\left(\mathbf{X}\right)$ and $D_{\mathbf{U}}^{2}f\left(\mathbf{X}\right)$. By the limit definition of the derivative, the first derivative of $f\left(\mathbf{X}\right)$ with respect to each entry of $\mathbf{X}$ can be expressed as follows:
$\frac{\partial f\left(\mathbf{X}\right)}{\partial X_{ij}}=\lim_{\Delta t\rightarrow 0}\frac{f\left(\mathbf{X}+\Delta t\mathbf{e}_{i}\mathbf{e}_{j}^{\top}\right)-f\left(\mathbf{X}\right)}{\Delta t},$ (7)
where $\mathbf{e}_{i}$ is the $i$th standard basis vector of $\mathbb{R}^{K}$ and $\mathbf{e}_{j}$ is the $j$th standard basis vector of $\mathbb{R}^{L}$. If the limit in (7) exists, then by substitution of variables
$\frac{\partial f\left(\mathbf{X}\right)}{\partial X_{ij}}U_{ij}=\lim_{\Delta t\rightarrow 0}\frac{f\left(\mathbf{X}+\Delta tU_{ij}\mathbf{e}_{i}\mathbf{e}_{j}^{\top}\right)-f\left(\mathbf{X}\right)}{\Delta t}.$ (8)
By definition, the total change in $f(\mathbf{X})$ in the direction of $\mathbf{U}$ is the sum of changes due to each entry of $\mathbf{X}$. Then
$D_{\mathbf{U}}f(\mathbf{X})$ $=\sum_{i,j\in[K]\times[L]}\frac{\partial f(\mathbf{X})}{\partial X_{ij}}U_{ij}$ (9)
$=\sum_{i,j\in[K]\times[L]}\lim_{\Delta t\rightarrow 0}\frac{f(\mathbf{X}+\Delta tU_{ij}\mathbf{e}_{i}\mathbf{e}_{j}^{\top})-f\left(\mathbf{X}\right)}{\Delta t}$ (10)
$=\lim_{\Delta t\rightarrow 0}\frac{f\left(\mathbf{X}+\Delta t\mathbf{U}\right)-f\left(\mathbf{X}\right)}{\Delta t}.$ (11)